OpenStudy (anonymous):

how do you factor equations? For example, 6x^2+24x=6?

OpenStudy (anonymous):

\[ 6 x^2 + 24 x -6=0\\ 6( x^2 + 4 x -1) \] Can you factor it more?

OpenStudy (anonymous):

You can if you want to use irrational numbers \[ 6 x^2+24 x-6=6 \left(x+\sqrt{5}+2\right) \left(x-\sqrt{5}+2\right) \]

OpenStudy (anonymous):

im not quite sure

OpenStudy (anonymous):

but without using irrational numbers, you wouldn't be able to right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

YW